OpenStudy (anonymous):

solve for x: 3/3x+1/x+4=10/7x

OpenStudy (shamil98):

Could you write this with the equation editor so we can see the exact problem you're asking?..

OpenStudy (anonymous):

\[\frac{ 3 }{ 3x }+ \frac{ 1 }{ x+4 }= \frac{ 10 }{ 7x }\]

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

\[\huge \frac{ 3 }{ 3x } + \frac{ 1 }{ x+4 } = \frac{ 10 }{ 7x }\] \[\huge \frac{ 1 }{ x } + \frac{ 1 }{ x+4 } = \frac{ 10 }{ 7x }\] \[\huge \frac{ 1(x+4) + 1(x) }{ x(x+4) } = \frac{ 10 }{ 7x }\] \[\huge \frac{ 2x+4 }{ x^2 + 4x } = \frac{ 10 }{ 7x }\] Cross multiply it now =)

OpenStudy (anonymous):

8x and 4x

OpenStudy (shamil98):

Cross multiply the terms... what do you mean 8x and 4x..

OpenStudy (shamil98):

\[\huge (2x+4)(7x) = 10(x^2 + 4x)\] Multiply, expand, get everything on one side.

OpenStudy (anonymous):

ohh i get for my answer x=-3

OpenStudy (anonymous):

@dalia_lam can u help?

OpenStudy (nikato):

u should have 2 answers again

OpenStudy (anonymous):

-3 and 3

OpenStudy (nikato):